OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Write the solution as an ordered pair

OpenStudy (anonymous):

X + y = 270 x - y = 26 Add the two equations: 2x = 296 x = 148 Substitute this value of x back into either of the original equations and solve for y: 148 + y = 270 y = 122 An ordered pair might be (148, 122)

OpenStudy (anonymous):

that was wrong\

OpenStudy (anonymous):

Exponents Quick Explanation: An exponent is a short-handed method of expression repeated multiplication. Rather than writing 5x5 we can simply write 52. They mean the same thing -- the superscript 2 means multiply five twice. Similarly, y4 means multiply y four times, i.e. y*y*y*y. More Detail: It doesn't seem all that hard to just write 5x5 instead of 52, but there are cases where the exponent could be quite large -- image writing out 520! Further down the road you'll also see that exponents can be negative, and don't even have to be whole numbers! Anyway, you'll learn that many, many functions in math and in nature have a squared dependence -- that means something depends on itself squared, or raised to the second power, or 52, or 5x5. They all mean the same thing. Believe it or not, exponents will make things easier... just have patience! How can I do math with exponents? If you understand that an exponent represents the number of times you multiply something, you can immediately understand what happens when we multiply two variables with exponents: Simplify this expression: x2*x6 Since x2 really just means x*x, and x6 just means 6 more x's, we end up with 8 x's multiplied together, right? Well what is an exponent -- the number of times we multiply something! Therefore, x2 times x6 equals x8. What did you learn from that example? When the same variable is multiplied, any exponents are added together. Adding to the exponent is the same as multiplying more times: Rule 1: (xa)*(xb) = xa+b There are other rules with exponents as well. If multiply two variables adds their exponents, then division must subtract exponents! Check out this example: Well, remember that this is just a quick way of writing this: Hopefully you remember enough basic algebra that you know to cancel a factor that's in the numerator and the denominator. In fact, we can scratch off three of the x's, leaving just the numerator: So that gives us another rule: Rule 2: Let's introduce a few more rules of exponents quickly: Rule 3: x1 = x That rules makes sense, because having just one x can't equal anything else but x, right? The next one might make a little less sense, but here it is: Rule 4: x0 = 1. (ANYTHING to the 0 power = 1). Here's why this is the case. With exponents, we are always dealing with multiplication. The identity in multiplication is 1. Imagine taking rule 1, and adding in a couple 0's: (xa)*(xb) = xa+b+0+0 Logically, we shouldn't have changed anything by adding those zeros. Since you can rewrite it as: xa*xb*x0*x0 In order for that to be the same as xa*xb, the x0 factors must equal 1.

OpenStudy (anonymous):

heres this to help u :)